# A look on equations describing pseudospherical surfaces

### v1

@@ 4 Final remarks
The investigation of PSS equations essentialy begun with the papers by Sasaki [38] and Chern and Tenenblat [8]. Both papers were motivated by the geometric nature of certain integrable equations in the sense describled in [1].
@@
The paper by Chern and Tenenblat [8], although motivated by Sasaki [38] started systematic geometric studies of integrable systems from the point of view of two-dimensional Riemannian manifolds. In consequence, nowadays we know that PSS equations are not only restricted to AKNS equations. In fact, equations with no integrable properties (in any sence) can be PSS equations, e.g., see [5] for some examples of non-integrable equations that are PSS equations.
@@
E. Reyes has made fundamental contributions in the study of PSS equations and integrable systems [3, 30, 31, 32, 33, 34, 35]. In particular, he proved that the CH equation is geometrically integrable [32]. In [33] the connections between PSS equations and integrability are explored, while in [35] a nice review on PSS in light of the legacy by Chern and Tenenblat contributions is presented.
@@
A topic not treated in this paper is the question of immersion, into the Euclidean space, of the abstract surfaces determined by the solutions of PSS equations. It is well known that PSS can be (locally) immersed into the three-dimensional space. In particular, there are works studying the immersion of surfaces determined by equations, see [6, 21, 24, 25, 26, 30] and references therein.
@@
Acknowledgements
I would like to thank Professor Tiago Macedo for his condition for me to contribute to CM with this review paper. I am also thankful to CNPq (grant nº 310074/2021-5) and FAPESP (grant nº 2024/01437-8) for financial support.

### v4

@@ 4 Further remarks
The investigation of PSS equations essentially began with the papers by Sasaki [38] and Chern and Tenenblat [5]. Both papers were motivated by the geometric nature of certain integrable equations in the sense described in [1].
@@
Chern and Tenenblat [5], although motivated by Sasaki [38], started systematic geometric studies of integrable systems from the point of view of two-dimensional Riemannian manifolds. Consequently, nowadays, we know that PSS equations are not only restricted to AKNS equations. In fact, equations without any integrable properties can be PSS equations, e.g., see [40] for some examples of non-integrable equations that are PSS equations.
@@
E. Reyes has made fundamental contributions to the study of PSS equations and integrable systems [3, 31, 32, 33, 35, 34, 36]. In [35] the connections between PSS equations and integrability are explored, while in [36] a nice review on PSS in light of the legacy of Chern and Tenenblat contributions is presented.
@@
Reyes proved that the CH equation is geometrically integrable [33], which is the starting point for the results established in [16], bringing qualitative information to the surface from the properties of the initial data of Cauchy problems involving the CH equation.
Remark 4.5.
A topic not treated in this paper is the question of immersion, into the Euclidean space, of the abstract surfaces determined by the solutions of PSS equations. It is well known that PSS can be (locally) immersed in the three-dimensional space. In particular, there are works studying the immersion of surfaces determined by equations, see [39, 18, 22, 23, 24, 31] and references therein.
@@
Remark 4.7.
From a topological viewpoint, all surfaces in Figure 1 possess singularities. On the other hand, the results established in [16, 18] made possible the “discovery” of a new kind of singularity: the blow-up of the metric. It is not clear to the authors whether the topological singularity of the surfaces in Figure 1 might be somewhat related to the blow up of the metric of the surfaces described by the solutions of the Camassa-Holm or Degasperis-Procesi equations experiencing wave-breaking. This is a point that should be better clarified in the future.
5 Discussion
The development of the theory of equations describing pseudospherical surfaces, as traced in this review, illustrates a profound and dynamic synergy between differential geometry and non-linear analysis. It was brought to light due to the elegant structure of the AKNS system [1], which played a vital role in the foundational works of Sasaki [38] and Chern–Tenenblat [5] that aimed to provide a geometric framework for the (at that time) emerging theory of integrable systems. Over time, however, the scope of the theory has expanded significantly. It is now widely recognized that the property of describing a pseudospherical surface is not confined to integrable equations. As shown by many works developed by Tenenblat and co-workers, e.g., see [39], certain differential equations possess this geometric attribute without admitting a Lax pair or soliton solutions, demonstrating that the class of PSS equations is larger than the class of integrable equations.
The contributions of Reyes [31, 32, 33, 35, 34, 36, 3] have been pivotal in refining the classification of PSS equations and clarifying the relationship between geometry and integrability, including extrinsic aspects [31, 3]. His introduction of the concept of geometric integrability provided a key distinction: it isolates those PSS equations whose associated pseudospherical structure depends essentially on a non-removable parameter. Such a parameter bridges the general geometric definition with the analytical machinery of the Inverse Scattering Transform, offering a geometric explanation for the presence of infinite hierarchies of conservation laws and Bäcklund transformations, which emerge from isospectral deformations of the surface.
More recently, a gap between the analysis of PDEs and geometry has been addressed by recent contributions of the author. A central issue addressed by recent research, and highlighted in this review, is the question of the regularity of solutions of PSS equations and their impact on the geometry determined by them. Classical formulations [5, 39, 33] were set assuming $C^{\infty}$ solutions, meaning that the differential forms are defined on infinite-order jet spaces—a natural assumption for classical soliton equations. However, the study of wave-breaking phenomena in models such as the CH equation requires moving beyond this smooth paradigm. In these models, solutions remain continuous while their derivatives become unbounded in finite time (breaking time), a behavior not captured by the infinite jets setting where the theory originally proposed by Chern and Tenenblat [5] was built. The results in [16, 18] show that the pseudospherical structure remains well-defined even for metrics of finite regularity, effectively transitioning the theory from infinite to finite jet spaces. This adaptation is not a contradiction but an essential evolution, enabling the geometric description of singular dynamics characteristic of peakon equations.
It is not yet clear to the author if the modifications introduced in Definition 2.1 and Definition 2.2 are sufficient for studying more complex and challenging solutions of equations like the CH. An example is the peakon solutions: given that they are distributional solutions of a non-local evolution for of the CH equation, they do not describe any surface in the sense of Definition 2.1. On the one hand, this shows an essential limitation of these recently introduced innovations. On the other hand, the requirement of finite jets in the definition of PSS tells us that we can expect issues when the problem of immersions is studied in conjunction with Cauchy problems, at least for CH type equations. This is a topic to be better understood in the field.
6 Conclusion
In this survey, we revisited the notions of PSS equations from its roots, based on the theory of integrable systems, passing through the fundamental contributions by Tenenblat and Reyes, until recent works of the author, who have incorporated tools of qualitative theory of PDEs in the study of PSS.